OpenStudy (anonymous):

Write the decimal as a fraction or a mixed number in the simplest form 3.4 with a line over the 4 A. 4 over 9 B. 3 4 over 9 C. 3 2 over 5 D. 3 4 over 11

jimthompson5910 (jim_thompson5910):

Let x = 3.444444... So 10x = 34.44444.... 10x - x = 34.44444.... - 3.44444.... 9x = 31 x = 31/9 From here, convert that last result into a mixed number.